For a p-electron, orbital angular moment is


(A) √2h                              (B)   h


(C) √6h                              (D)   2h                    plz EXPLAIN  IT PROPER....

Grade:Upto college level

2 Answers

aditya reddy
13 years ago

O.A.M= SQRT[ l*(l+1)]   * h

 

FOR P ELECTRON  L VALUE IS 1.

 

THEREFORE O.A.M= SQRT(2) * h             

 

THEREFORE OPTION---A

deeksha sharma
13 years ago

orbital angular momentum of an orbital is given by {l(l+1)}1/2  h


for p orbital, l=1          (s=0

                                 p=1

                                 d=2

                                 f=3)

o angular momentum= {1(1+1)}1/2 h ={1*2}1/2   h  = 21/2 h

so the answer is A under root pf 2 multiplied by h